OpenStudy (anonymous):

find the vertices and foci of the hyperbola: Y^2/5-X^2=1

OpenStudy (anonymous):

The foci of this hyperbola are found using the Pythagorean Theorem \[a^2+b^2=c^2\] Whereas the vertices of the parabola are given as \[(0,\pm a)\] The equation of this hyperbola follows the form \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1 \] Can you solve the problem now?

OpenStudy (anonymous):

not really i need further explaining

OpenStudy (anonymous):

What didn't you understand?

OpenStudy (anonymous):

what # to put where

OpenStudy (anonymous):

OK, so the equation is \[\frac{y^2}{5}-x^2=1\] equivalent to \[\frac{y^2}{5}-\frac{x^2}{1}=1\] so \[a^2=5, b^2=1\]